how do I solve x to the -3 power in parenthesis and a 2 outside the parenthesis? I know for the negative exponent I make it 1/x to the 3rd but how do I square a fraction?
Answer by stanbon(48535)

how do I solve x to the -3 power in parenthesis and a 2 outside the parenthesis? I know for the negative exponent I make it 1/x to the 3rd but how do I square a fraction?
-----------
(x^-3)^2
(1/x^3)^2
= (1/x^3)*(1/x^3) = 1/x^6
==============================
Cheers,
Stan H.